### v2

# What LLMs Know but Don’t Say: Non-generative Prior Extraction for Generalization

Abstract
In domains like medicine and finance, large-scale labeled data is costly and often unavailable, leading to models trained on small datasets that struggle to generalize to real-world populations. Large language models contain extensive knowledge from years of research across these domains. We propose LoID (Logit-Informed Distributions), a deterministic method for extracting informative prior distributions for Bayesian logistic and linear regression by directly accessing their token-level predictions. Rather than relying on generated text, we probe the model’s confidence in opposing semantic directions (positive vs. negative impact) through carefully constructed sentences. By measuring how consistently the LLM favors one direction across diverse phrasings, we extract the strength and reliability of the model’s belief about each feature’s influence. We evaluate LoID on 15 real-world tabular datasets under synthetic out-of-distribution (OOD) settings characterized by covariate shift. We compare our approach against (1) standard uninformative priors, (2) AutoElicit, a method that prompts LLMs to generate expert-style priors, (3) LLMProcesses, a method that prompts LLMs to generate predictions and (4) an oracle-style upper bound derived from fitting logistic/linear regression on an in-distribution (IID) train set. We assess performance using Area Under the Curve (AUC) and Mean Squared Error (MSE). Across datasets, LoID significantly improves performance over logistic and linear regression models trained on OOD data, recovering up to 50% of the performance gap relative to the oracle model.
1 Introduction
Machine learning models used in critical domains like medicine and finance often lack access to large, well-balanced, and comprehensively labeled datasets. Clinical datasets may cover only a single hospital or demographic, while credit-risk tables can be skewed toward historical lending patterns. For example, Purushotham et al. (2017) benchmarked several deep learning models on clinical prediction tasks—including mortality and length of stay—using the MIMIC-III intensive care electronic health record dataset from a single hospital site. As a consequence, models trained on small, non-representative samples frequently fail to generalize when confronted with real-world populations that differ from their training distribution Brigato and Iocchi (2020); Li et al. (2024); Althnian et al. (2021). Bayesian methods offer a principled way to encode prior knowledge and mitigate such brittleness, but manually eliciting high-quality priors at scale remains a longstanding challenge Mikkola et al. (2023).
LLMs, trained on vast corpora of scientific literature and domain-specific data Brown et al. (2020), offer a promising new source of priors for downstream modeling. Recent efforts have explored LLMs both as zero-shot and few-shot predictors for tabular classification Hegselmann et al. (2023); Requeima et al. (2024), and as in-context learners capable of producing joint probabilistic predictions Shysheya et al. (2025). Other work has investigated the integration of LLMs with classical statistical methods: LLM-FE Abhyankar et al. (2025) uses LLMs to guide feature engineering, LLM-Lasso Zhang et al. (2025) for feature selection, and LLMCorr Zhong et al. (2024) for post-hoc model correction. AutoElicit Capstick et al. (2025) introduced a technique to query LLMs for Gaussian priors, but its reliance on generative sampling introduces run-to-run variability.
While LLMs offer a promising source of priors for downstream modeling, their practical benefit depends on how well these priors generalize under distribution shift. OOD generalization has therefore received increasing attention, as real-world deployment often involves data that differs from the development environment. Meta-learning approaches like TabPFN Hollmann et al. (2023) and Drift-Resilient TabPFN Helli et al. (2024) address OOD robustness through synthetic task generation and causal modeling. Others have proposed logit-based calibration strategies such as UE-NL Huang and Qiao (2023), or LLM-guided data augmentation methods like Summary Boosting Manikandan et al. (2023) and LLMOverTab Isomura et al. (2025) to cover underrepresented regions of the feature space. However, despite this momentum, few works have systematically explored how to extract priors from the internal structure of LLMs (e.g. hidden states, logits, etc.) for tabular tasks under distribution shift.
We address this gap with LoID (Logit-Informed Distributions), a deterministic framework that elicits informative priors for Bayesian logistic and linear regression directly from an LLM’s token-level logits. Instead of relying on generated completions or examples, LoID probes the LLM with semantically paired sentences highlighting positive and negative feature-target relationships. The logit difference between the relational tokens quantifies directional belief, and the variance across paraphrased prompts yields an uncertainty estimate—together forming Normal priors over model coefficients. These priors integrate into Bayesian workflows, without any fine-tuning or in-context learning.
We evaluate LoID on 14 publicly available tabular datasets that are widely used benchmarks and have been studied in prior work, along with one non-public medical dataset to assess performance on data that does not exist in pre-training corpora. We compare LoID against uninformative Bayesian inference, LLMProcesses, and AutoElicit. To provide an upper bound, we additionally train logistic and linear regression models on IID training sets, offering insight into performance when the true data distribution is available.
Our experiments show that LoID improves OOD generalization across 11/15 tabular datasets, recovering 50% of the lost performance on average.
In summary, our contributions include:
  • Logit-based deterministic prior elicitation. LoID is the first method to extract Bayesian priors directly from the token-level logits of LLMs, eliminating the variability associated with text generation and prompt sampling.
  • Baseline comparison across 15 OOD datasets. We evaluate LoID on diverse OOD tabular datasets from healthcare, finance, and social science. Our experiments demonstrate that LoID often outperforms existing approaches across most settings.
2 Related work
2.1 LLMs for tabular data modeling
The integration of LLMs into tabular data modeling has garnered attention, particularly in addressing challenges such as limited data availability, complex feature interactions, and the necessity for domain-specific feature engineering.
Several studies have explored the use of LLMs as standalone predictors for tabular data. Hegselmann et al. (2023) introduce TabLLM, a framework that applies large language models to zero-shot and few-shot classification of tabular data by converting rows into natural-language strings, enabling effective classification even with minimal labeled examples. Requeima et al. (2024) propose LLMProcesses, a framework that enables LLMs to produce probabilistic regression outputs by conditioning on both numerical data and natural language descriptions of prior knowledge. Shysheya et al. (2025) present JOLT, a method that leverages the in-context learning capabilities of LLMs to define joint probabilistic predictions over tabular data.
Beyond standalone applications, recent research has focused on leveraging LLMs to enhance classical statistical models. For instance, Abhyankar et al. (2025) introduced LLM-FE, a framework that combines evolutionary search with the reasoning capabilities of LLMs to iteratively generate and refine feature transformations for tabular data. Similarly, LLM-Lasso Zhang et al. (2025) is a domain-informed approach to feature selection and regularization by leveraging large language models to guide the selection process. Capstick et al. (2025) proposed AutoElicit, a method that interacts with LLMs to extract priors for Gaussian-based probabilistic models. In a different vein, Zhong et al. (2024) presented LLMCorr, a training-free framework where LLMs act as post-hoc correctors to propose corrections for the predictions of arbitrary machine learning models. Liu et al. (2023) investigated the reasoning capabilities of LLMs over tabular data and proposed combining symbolic and textual reasoning to improve performance on table-based question answering tasks.
2.2 Out-of-distribution generalization
OOD generalization is a fundamental challenge in real-world applications of tabular machine learning. Models trained on limited or biased datasets often fail to generalize when deployed on test data that differ from the training distribution. A variety of approaches have been proposed to address this issue through data augmentation, distributionally robust optimization, causal modeling, Bayesian inference, and most recently, the use of LLMs.
One line of work focuses on meta-learning for small-data generalization. Hollmann et al. (2023) introduced TabPFN, a foundation model trained on millions of synthetic tabular tasks to approximate Bayesian posterior predictive distributions in one forward pass. The model shows strong generalization across novel tasks and distributions. Building on this, Helli et al. (2024) proposed Drift-Resilient TabPFN, which introduces temporally shifting causal mechanisms during training to improve robustness under distribution shift.
Several recent studies have explored logit-based and uncertainty-aware methods for detecting or adapting to OOD inputs. Huang and Qiao (2023) proposed Uncertainty Estimation with Normalized Logits (UE-NL), a calibration technique that regularizes softmax confidence and improves both in-distribution accuracy and OOD detection.
In the LLM context, OOD robustness has been approached in several creative ways. Summary Boosting (Manikandan et al., 2023) uses GPT-generated feature summaries as weak learners to improve robustness in small or under-specified datasets. LLMOverTab (Isomura et al., 2025) performs data augmentation by prompting LLMs to generate realistic examples, improving classifier performance on underrepresented regions of the feature space. These strategies demonstrate that LLMs can inject external knowledge or expand the training distribution to mitigate generalization failure.
3 Method
We propose LoID (Logit-Informed Distribution), a deterministic framework for extracting priors for tabular dataset features from LLMs using token-level logits, rather than relying on generated samples. Our goal is to derive informative priors for Bayesian logistic and linear regression that generalize robustly in low-data settings.
3.1 Problem setting
Let $\mathcal{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote a training dataset, where each $x_{i}\in\mathbb{R}^{d}$ is a vector of $d$ features and $y_{i}\in\{0,1,\ldots,m\}$ is the corresponding label. In Bayesian logistic regression, we model the likelihood as:
$p(y_{i}=t\mid x_{i},\boldsymbol{\beta})=\sigma(x_{i}^{\top}\boldsymbol{\beta}),$
where $\sigma(\cdot)$ denotes the sigmoid function and $\boldsymbol{\beta}\in\mathbb{R}^{d}$ represents the regression coefficients. Bayesian inference is performed by specifying a prior distribution $p(\boldsymbol{\beta})$ and computing the posterior $p(\boldsymbol{\beta}\mid\mathcal{D}_{\text{train}}).$ In data-scarce settings, informative priors can substantially improve generalization. LoID constructs such priors by leveraging domain knowledge encoded in LLMs, without access to the training data.
3.2 Querying LLM beliefs via logit differences
For each feature $f_{j}$, we construct paired prompts that express opposing semantic claims about its effect on the target variable. Each prompt is carefully designed such that all relevant context appears before a sentiment-bearing token (e.g., “positive” or “negative”), ensuring that the token’s probability considers the entire context for the task, including the name of the feature and the prediction task, $t$ (see examples in A.2.1). For example, given a prompt prefix like:
‘‘The impact of $f_{j}$ on $t$ is ’’
We use the probabilities from the LLM’s softmax over its vocabulary to extract the probability of the next token being positive or negative. For the $i$th prompt prefix we define:
$\mathcal{P}_{i}^{+}=P(\texttt{<positive\_token>}\mid\text{prefix}),\quad\mathcal{P}_{i}^{-}=P(\texttt{<negative>\_token}\mid\text{prefix})$
We then define the logit preference score as:
$\text{logit}_{j}(p_{i})=\log\left(\frac{p_{i}}{1-p_{i}}\right),\quad p_{i}=\frac{P_{i}^{+}}{P_{i}^{+}+P_{i}^{-}}$
The logit preference score quantifies the extent to which the model prefers the positive token over the negative token in a given context. While this difference may also reflect other patterns learned during training, our hypothesis is that by constructing sentences that do not grammatically or syntactically favor either token, we reduce the influence of such confounding factors. Finally, applying the logit transformation maps probabilities from the bounded interval (0, 1) to the unbounded real line, making the resulting scores directly appliable as coefficients in logistic regression models.
Figure 1: LoID Pipeline: To extract informative priors, LoID uses multiple sentence templates for each feature and computes the probabilities of the tokens "positive" and "negative" as the next token. From these probabilities, we calculate the logit preference score, which quantifies the model’s preference for one token over the other. Using the logit preference scores from all sentences, we compute their mean and standard deviation. We then apply a transformation to the standard deviation to reduce overconfidence, allowing the data to have more influence on the resulting distribution.
3.3 Constructing priors from logits
We model each coefficient $\beta_{j}$ with a Gaussian prior: $\beta_{j}\sim\mathcal{N}(\mu_{j},\sigma_{j}^{2})$. The mean for feature $f_{j}$ is given by averaging this score across a set of paraphrased prompts T{T}:
$\mu_{j}=\frac{1}{|T|}\sum_{i\in T}\text{logit}_{j}(p_{i})$
This value reflects the LLM’s normalized preference for a positive versus negative association with the feature. Figure 1 demonstrates how LoID extracts informative priors from model logits. To capture uncertainty in the LLM’s belief, we define the standard deviation $\sigma_{j}$ as:
$\sigma_{j}^{2}=\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
Here, $logit_{i}(p_{j})$ denotes the logit preference score for the $i$th prompt and feature $j$. This design assigns higher variance to features for which the LLM shows inconsistent preferences across paraphrases, capturing greater uncertainty about their influence. We evaluated three variance computation methods and concluded that taking the standard deviation of logit preference scores as described above performs best (Section A.2.3).
3.4 Bayesian inference with LoID priors
With LoID priors, posterior inference is performed using a Laplace approximation, which approximates the posterior with a Gaussian centered at the maximum a posteriori (MAP) estimate:
$\hat{\beta}=\arg\max_{\beta}\log p(\beta\mid D_{\text{train}})$ (1)
The posterior is then locally approximated via a second-order Taylor expansion of the log-posterior:
$p(\beta\mid D_{\text{train}})\approx\mathcal{N}(\hat{\beta},\Sigma),\quad\Sigma=\left[\nabla^{2}_{\beta}\left(-\log p(\beta\mid D_{\text{train}})\right)\big|_{\beta=\hat{\beta}}\right]^{-1}$ (2)
This provides efficient uncertainty estimates without sampling, with LoID priors affecting both the MAP location and posterior curvature. All results are computed under this Gaussian approximation.
3.5 Evaluation
To evaluate LoID, we construct synthetic OOD splits from real-world datasets. We begin by creating a standard 80–20 random split, where the training portion serves as the in-distribution (IID) training set and the held-out portion serves as the IID test set. To simulate distribution shift, we derive an OOD training set by selecting a subset of the IID training data and systematically excluding specific categories of categorical features or ranges of numerical features. Section A.5 explains the synthetic distribution shift for each dataset in detail.
We verify the presence of distribution shift by comparing the predictive performance of a model trained on the IID training set with that of a model trained on the OOD training set, both evaluated on the same IID test set. A substantial degradation in performance when training on the OOD split indicates that the training data no longer adequately represents the true data distribution.
Given this setup, we then train Bayesian models with LoID priors using the OOD training set and evaluate their performance on the IID test set. Our primary objective is to assess the extent to which LoID priors can mitigate the performance gap introduced by distribution shift.
4 Experimental setup
4.1 Datasets
Head and Neck Cancer Recurrence dataset
We compiled a binary classification dataset for Head and Neck Cancer Recurrence. This dataset includes comprehensive features about patient demographics, tumor characteristics, treatment regimens, and clinical staging, and predicts the likelihood of cancer recurrence following radiotherapy. The dataset encompasses 508 patients with head and neck cancers, capturing clinical variables including comorbidity indices (Charlson score), smoking history (pack-years), HPV status, primary tumor site localization, and detailed radiation therapy parameters such as treatment duration, prescription dose, and the critical time interval between biopsy and treatment initiation.
This dataset presents a significant class imbalance (390 non-recurrence vs. 118 recurrence cases), reflecting real-world clinical distributions. The dataset is particularly valuable for evaluating model performance under distributional shift, as we created an out-of-distribution test set focusing on elderly patients ($age\geq 75$). This datasets represents a clinically meaningful population that is not present in any LLM pre-training corpora. [^1] We will release this dataset upon publication.
Public datasets
We evaluate LoID on 14 publicly available tabular datasets for binary classification and regression: Adult Income Becker and Kohavi (1996), Cervical Cancer Fernandes et al. (2017), Chronic Kidney Disease Rubini et al. (2015), Bank Marketing Moro et al. (2014) Blood Transfusion Service Center Yeh (2008), Stroke Fedesoriano (2020), Heart Failure UCI Machine Learning Repository (2020), Diabetes Prediction Mustafatz (2023), Give Me Some Credit Kaggle (2011), Airfoil Self-Noise, Bike Sharing Fanaee-T (2013), Combined Cycle Power Plant Tfekci and Kaya (2014), Concrete Compressive Strength, Yeh (1998), and Communities and Crime Redmond (2002).
Preprocessing
Across all datasets, we apply the same preprocessing pipeline: categorical features are one-hot encoded, missing categorical values are imputed with mode and missing numerical values are imputed with mean. For datasets with an excessive number of features—namely Bank Marketing, Breast Cancer Diagnostic, Cervical Cancer, Chronic Kidney Disease, Communities and Crime, Heart Failure, Haptitis, and Head and Neck Cancer Recurrence—we perform feature selection using Claude-3.5-Sonnet by providing feature descriptions and prompting it to return a list of the 15 most relevant features for the prediction task. The same reduced feature set is used consistently across all experimental conditions to ensure comparability. We apply z-score normalization to the features for AutoElicit and LoID, but not for LLMProcesses, following the authors’ recommendations.
Development datasets
We designate the Stroke, Blood Donation, Heart Failure, and Adult Income datasets as development datasets for choosing an LLM and hyperparameter tuning A.2. The remaining 11 datasets are reserved for testing the generalization performance of the final LoID method.
4.2 Creating distribution shifts
For each dataset, we construct synthetic OOD splits as described in Section A.5. As shown in Table 1 in the P_IID (in-distribution train set performance), and P_OOD (out-of-distribution train set performance) columns, the model trained on the OOD train set noticeably underperforms compared to the model trained on the IID train set. This confirms that generalization is impaired and the OOD train set does not represent the full distribution of the dataset as well as the IID train set.
Dataset P_OOD P_IID N(0,1) LoID LLMProc AutoEl Gap (%)
Cap our method Requeima et al. (2024) Capstick et al. (2025)
Classification (AUC)
Head & Neck Cancer 0.542 0.998 0.567 0.696 0.540 0.609 33.8
Adult Income 0.414 0.809 0.51 0.745 0.603 0.510 83.8
Cervical Cancer 0.440 0.870 0.448 0.681 0.658 0.570 56.0
Kidney Disease 0.822 0.993 0.715 0.901 0.590 0.320 46.2
Diabetes 0.883 0.963 0.853 0.906 0.502 0.864 28.8
Stroke 0.608 0.842 0.618 0.694 0.500 0.670 36.5
Heart Failure 0.746 0.849 0.742 0.753 0.770 0.748 6.8
Bank Marketing 0.780 0.878 0.780 0.652 0.498 0.692 -130.6
Blood Donation 0.572 0.748 0.572 0.663 0.483 0.556 51.7
Good Credit Score 0.710 0.790 0.710 0.760 0.700 0.710 62.5
Regression (MSE)
Concrete Strength 469.20 98.14 165.72 120.59 136.57 1974.11 93.9
Communities & Crime 0.17 0.01 0.073 0.07 0.15 0.30 62.5
Airfoil Noise 77.16 23.69 51.11 60.51 47.80 62.21 68.9
Cycle Power 962.94 21.24 23.58 509.92 462.89 47.14 0.1
Bike Sharing 77858 19030 71890 78351 75650 71405 -1.8
Table 1: Performance comparison across 15 datasets. Gap Closed represents the percentage of the OOD performance gap recovered by our method, LoID. The percentage closed in calculated using: $\frac{|LoID-P\_OOD|}{|P\_IID-P\_OOD|}\times 100$
4.3 LLM-based prior extraction
We use Gemma-2-27B, Gemma-3-27B, llama-3-8B, and GPT-2 to extract logit-based priors as described in Section 3.2. We use these models in zero-shot mode without any fine-tuning or in-context examples. In our experiments, the queried lexical items are represented as single tokens across models, enabling direct logit-based comparisons.
Each dataset is processed independently. For each feature, we issue 10 prompts to extract token-level log-probabilities for the positive and negative completions. The number of prompts is selected via hyperparameter tuning (Section A.2.1). The resulting feature-wise prior means and variances are then used to define informative Gaussian priors in Bayesian logistic and linear regression models. Posterior inference is performed using PyMC with a Laplace approximation.
We compare LoID to the following baselines.
  • P_OOD: Logistic/Linear regression model trained on OOD train set.
  • AutoEl (AutoElicit): LLM-generated expert-style priors via sampled completions.
  • LLMProc (LLMProcesses): In-context learning approach that generates numerical predictions by formatting tabular data as natural language prompts. For fair comparison, we use the OOD training examples as few shot examples for the LLM.
  • P_IID: Performance obtained by training a logistic/linear regression model on the in-distribution (IID) data. This represents the oracle benchmark where the model has full access to the training distribution, serving as a reference point for evaluating OOD generalization.
  • N(0,1): An uninformed baseline using standard normal $\mathcal{N}(0,1)$ priors for all feature coefficients. This represents a weakly informative Bayesian approach without incorporating domain knowledge, serving as a reference point to quantify the information gain provided by LLM-derived priors. While a uniform prior would be maximally uninformative, our experiments in Section A.2.4 show that normal priors are more compatible with the MCMC pipeline and provide a stricter baseline.
For AutoElicit and LLMProcesses, we adopt the hyperparameters recommended by the original authors to the extent possible in order to ensure a fair comparison. A comprehensive list of the hyperparameters used to reproduce these baselines is provided in Section A.1.
5 Results and discussion
We evaluate LoID across 15 datasets spanning diverse domains to understand when LoID is able to close the performance gap caused by covariate shift. Our results in Table 1 reveal a systematic pattern: LoID excels when feature-target relationships are semantically meaningful and generalizable, but struggles when relationships are contextual, behavioral, or dataset-specific.
5.1 When does LoID help?
Medical and demographic tasks with stable relationships (83.8% gap closed on average). LoID achieves the strongest gains on datasets where feature-target relationships reflect generalizable biomedical or socioeconomic knowledge:
  • Adult Income (83.8% gap): Age, education, and occupation have well-known effects on income that generalize across demographic shifts
  • Concrete Strength (93.9% gap): Physical properties (cement, age, water) have deterministic relationships with compressive strength
  • Chronic Kidney Disease (46.2% gap): Clinical markers (hemoglobin, albumin) have established diagnostic value
In these domains, the LLM’s world knowledge captures stable causal mechanisms that remain valid under distribution shift.
Physical systems with known causal structure. For regression tasks where features have clear physical interpretations (Concrete Strength: 93.9% gap, Airfoil Noise: 68.9% gap, Communities and Crime: 62.5% gap), LoID leverages the LLM’s understanding of domain physics and social science to regularize predictions under shift.
Figure 2: LLM prior alignment across success and failure modes. Top: Well-aligned priors for Blood Donation and Chronic Kidney Disease show strong agreement between LLM beliefs and actual correlations. Bottom: Combined Cycle Power exhibits directionally correct but magnitude-misspecified priors, while Bank Marketing shows absent priors (near-zero) for context-dependent features.
5.2 When does LoID hurt?
Behavioral and contextual tasks with dataset-specific patterns. LoID fails on three datasets where relationships are inherently local or behavioral:
  • Bank Marketing ($130.6\%$ gap): Customer banking behavior (call duration, previous contacts) varies wildly across campaigns.
  • Bike Sharing ($1.8\%$ gap): Demand is driven by local weather, events, and infrastructure, relationships that are city-specific and temporal
  • Combined Cycle Power ($0.1\%$ gap): Plant efficiency depends on equipment-specific calibrations and operational conditions that LLMs cannot infer from generic turbine descriptions
In these cases, the LLM’s priors encode generic relationships (e.g., “higher temperature decreases power output”) that may be directionally correct but quantitatively misspecified for the specific plant, city, or campaign. This introduces harmful bias that degrades performance.
Understanding failure modes through prior analysis. Examining the extracted priors reveals two distinct failure modes:
  1. 1.
    Overconfident but wrong priors destroy beneficial regularization. For Combined Cycle Power, the uninformed $\mathcal{N}(0,1)$ baseline closes 99.8% of the OOD gap (MSE: 23.58 vs. P_OOD: 962.94), demonstrating that simple Bayesian regularization is highly effective. However, LoID closes only 0.2% (MSE: 961.04), because the LLM’s equipment-agnostic priors (e.g., “higher ambient temperature decreases turbine efficiency”) override the beneficial regularization with domain knowledge that’s directionally plausible but quantitatively misspecified for this specific power plant and severe temperature shift (train: 1.8–13.5°C, test: 13.5–37.1°C). The LLM knows generic thermodynamics but cannot account for plant-specific calibrations or extreme extrapolation regimes.
  2. 2.
    Context-dependent features produce uninformative priors. For Bank Marketing and Bike Sharing, LLMs have no basis for judging how campaign-specific variables (e.g., “days since last contact”) or city-specific infrastructure affects outcomes. The extracted priors are either near-zero (indicating maximal uncertainty) or directionally inconsistent, providing no signal over uninformed regularization. Unlike Combined Cycle Power, where the problem is wrong priors, here the problem is absent priors—the LLM simply lacks relevant knowledge about these context-dependent domains.
Figure 2 illustrates four representative cases that reveal when and why LLM priors succeed or fail.
5.3 Comparison to baselines
LoID vs. LLMProcesses. LLMProcesses attempts to directly predict outcomes by prompting the LLM with feature values. While this works for more common tasks, it fails on real-world data: LLMProcesses achieves near-random performance (0.500 AUC) on 4/9 classification tasks (Stroke, Diabetes, Heart Failure, HNC). LoID outperforms LLMProcesses on 12/14 datasets (excluding Bank Marketing where both fail). This demonstrates that extracting priors is more effective than direct prediction—the LLM knows the relationships but cannot directly predict outcomes from sparse feature descriptions.
LoID vs. AutoElicit. AutoElicit elicits structured expert priors by prompting the LLM multiple times and aggregating responses to reduce variability. It performs well on tasks involving physical systems (Bike Sharing: 71405 MSE; Combined Cycle Power: 47.14 MSE), where aggregating numeric estimates helps calibrate prior ranges. However, for tasks requiring semantic reasoning about feature importance (Adult Income: 0.510 AUC; Concrete Strength: 1974.11 MSE), free-form elicitation may struggle to capture directional relationships as effectively as token-level probability extraction.
Moreover, LLMs tend to produce overconfident priors when explicitly prompted, which can limit the model’s ability to adapt to data. In contrast, LoID derives priors directly from token probabilities, leading to more calibrated uncertainty. As a result, LoID outperforms AutoElicit on 12/14 datasets.
The N(0,1) baseline reveals when priors help. Comparing LoID to the uninformed $\mathcal{N}(0,1)$ baseline isolates the value of LLM knowledge. On 11/15 tasks, LoID substantially outperforms $\mathcal{N}(0,1)$, confirming that LLM priors are informative.
6 Limitations and future work
While our results demonstrate the effectiveness of LoID, several limitations should be noted, which also point toward directions for future work.
  • LoID is most useful for dataset with meaningful semantic features, limiting its usage for de-identified datasets.
  • LoID requires access to token-level logits, limiting compatibility with closed-source APIs.
7 Conclusion
We introduced LoID, a method for extracting Bayesian priors from large language models through token-level probability distributions. By probing LLMs for beliefs about feature–target relationships and converting these into structured priors, LoID leverages embedded world knowledge to improve predictive performance under distribution shift.
Across 15 datasets, we observe a consistent pattern: LoID performs strongly when feature–target relationships are semantically stable and generalizable, closing 83.8% of the performance gap induced by out-of-distribution training in medical and demographic settings, but degrades when these relationships are context-dependent or dataset-specific, as in behavioral tasks such as Bank Marketing. Overall, LoID demonstrates that extracting distributional priors is more robust than direct prediction (compared to LLMProcesses) and more reliable than free-form elicitation (compared to AutoElicit), outperforming both baselines on 12 out of 15 datasets.
Appendix A Appendix
A.1 Baseline hyperparameters
This section details the hyperparameters used for the baseline methods compared in our experiments. Table 2 provides a comprehensive overview of all configuration settings.
Category LLMProcesses AutoElicit LoID
Model Configuration
Model Gemma-2-27B-it Gemma-2-27B-it Gemma-2-27B-it
Temperature 1.0 1.0 N/A (no sampling)
Top-p 1.0 0.9 N/A
Max length 20 tokens 8192 tokens N/A
Prior/Prediction Extraction
Method Monte Carlo sampling Multi-expert prompting Token probability
Samples per point 50 500 (prior predictive) 10 templates
Aggregation Median MCMC mixture Mean and Variance of log-probs
Templates/Roles N/A 100 (10×10 role pairs) 10 sentence templates
Data Configuration
Training samples 50 (30 for HNC) 25 N/A (no in-context)
Prompt ordering Distance-based N/A N/A
Reproducibility
Random seed N/A 42 42
Repetitions 5 5 5
Table 2: Hyperparameter configurations for AutoElicit, LLMProcesses, and LoID.
A.2 Hyperparameter optimization
We used four datasets—Stroke, Blood Donation, Hear Failure, and Adult Income—as development datasets to tune hyperparameters based on AUC and MSE.
A.2.1 Number of prompt templates
We tested our method with 2–20 prompt templates and analyzed how the number of templates affects LoID’s performance. The following are some examples of our prompt templates:
  1. 1.
    The impact of on is …
  2. 2.
    The relationship between and is …
  3. 3.
    The role of in is …
  4. 4.
    When considering , the effect on is …
  5. 5.
    The correlation between and is …
Figure 3 reveals the optimal template count for prior extraction. Performance peaks at 10 templates across most datasets, balancing signal aggregation against overfitting. Beyond 10 templates, performance degrades sharply (e.g., Stroke drops from 0.69 to 0.39 at 15 templates), suggesting that excessive prompting amplifies misspecified priors rather than refining them. While coefficient of variation decreases with more templates (indicating more stable uncertainty estimates), this stability comes at the cost of performance when templates exceed 10. Notably, heterogeneity reduction plateaus after 10 templates, yet performance continues to degrade, demonstrating that template averaging helps only when the underlying priors are well-specified. For datasets where LLM knowledge is weak (e.g., Blood), even 2 templates outperform 15+, as additional prompting merely reinforces uninformative/incorrect beliefs. This pattern supports our recommendation of 10 templates as the optimal trade-off between variance reduction and avoiding overconfident misspecification.
Figure 3: Template count sensitivity analysis across four datasets. Top two rows: Performance (Test AUC) versus number of sentence templates, comparing LoID to out-of-distribution baseline (P_OOD). Performance peaks at 10 templates for most datasets, with degradation at 15+ templates suggesting overfitting to misspecified priors. Bottom two rows: Prior heterogeneity measured by coefficient of variation (CV) of extracted prior standard deviations. Higher CV indicates greater variance in uncertainty across features. Heterogeneity decreases with more templates as averaging stabilizes estimates, but excessive templates (15+) harm performance without further reducing uncertainty.
A.2.2 Model selection
We compared the performances of LLMs–GPT-2 (GPT), LLaMA-3.1-8B (L-3), Gemma-3-27B-Instruct (G-3), and Gemma-2-27B-Instruct (G-2) in terms of AUC. The best average performance was achieved with Gemma-2-27B, as shown in Table 3.
Dataset G-2 G-3 L-3 GPT-2
Stroke 0.694 0.690 0.624 0.712
Blood 0.663 0.522 0.609 0.579
Heart Failure 0.753 0.767 0.737 0.585
Adult 0.745 0.562 0.550 0.370
Average 0.714 0.635 0.630 0.562
(a) Model ablation.
Dataset M-1 M-2 M-3
Stroke 0.694 0.686 0.690
Blood 0.663 0.576 0.575
Heart Failure 0.753 0.708 0.743
Adult 0.745 0.674 0.659
Average 0.714 0.661 0.667
(b) Variance method ablation using Gemma-2-27B.
Table 3: Choosing the best model and variance estimate using four development datasets (Stroke, Blood Donation, Heart Failure, Adult Income).
A.2.3 Variance estimation formula
The prior standard deviation $\sigma_{j}$ for feature $j$ quantifies the uncertainty in the elicited prior mean. We explored two methods for computing $\sigma_{j}$ from LLM token probabilities, each capturing different aspects of elicitation uncertainty. The best average performance was achieved with Method 1, as shown in Table 3.
Method 1: variance of normalized logit probabilities
In this method, $p_{j}$ reflects the model’s preference between logits. Applying a log transform maps probabilities from $[0,1]$ to the real line, aligning with the unbounded nature of regression coefficients. Variance across paraphrased inputs captures inconsistency in LLM responses, indicating greater epistemic uncertainty.
$\sigma_{j}^{2}=\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
where
$\text{logit}(p_{j})=\log\left(\frac{p_{j}}{1-p_{j}}\right)$
and
$p_{j}=\frac{P_{j}^{+}}{P_{j}^{+}+P_{j}^{-}}$
Method 2: entropy-based uncertainty
This method uses Shannon entropy to measure the uncertainty in the LLM’s token probability distribution. Higher entropy (closer to $\log 2\approx 0.693$) indicates the LLM is uncertain between the positive and negative tokens, reflecting ambiguity about the feature’s effect direction. Lower entropy indicates confident predictions.
$\sigma_{j}=\frac{1}{N_{\text{sent}}}\sum_{i=1}^{N_{\text{sent}}}H({P_{j}^{-}},P_{j}^{+})$
Method 3: standard deviation of log-probability differences
This method uses the difference of log-probabilities (equivalent to the log-ratio $\log(\frac{p_{\text{pos}}}{p_{\text{neg}}})$) without normalization, preserving information about the absolute probability mass assigned to both tokens. When templates produce inconsistent log-probability differences, this indicates epistemic uncertainty about both the direction and confidence of the feature’s effect. Unlike the normalized logit approach, this method is sensitive to cases where the LLM assigns low absolute probability to both tokens (indicating high uncertainty) versus high probability to one token (indicating confidence).
$\sigma_{j}^{2}=\text{std}(diff_{1}(p_{j}),\ldots,diff_{|T|}(p_{j}))$
Where
$diff(p_{j})=[\log(P_{j}^{+})-\log(P_{j}^{-})]$
A.2.4 Uninformative baseline: $\mathcal{N}$(0, 1) vs. $\mathcal{U}$(-1, 1)
We include an additional $\mathcal{N}(0,1)$ baseline to illustrate the improvement achieved over a more weakly informative prior that is commonly used with Bayesian inference.
To understand the impact of prior distribution choice, we compare two configurations: (1) uninformative Normal priors $\mathcal{N}(0,1)$, and (2) uninformative Uniform priors $\mathcal{U}(-1,1)$.
Table 4 reveals that while Uniform priors ostensibly provide “maximum uninformativeness” by assigning equal probability across a range, they fail as regularizers for regression coefficients because:
  1. 1.
    Lack of soft regularization: Uniform distributions do not encode the reasonable prior belief that smaller coefficient magnitudes are more plausible than larger ones. This absence of gradient-based regularization toward zero can lead to poor MCMC sampling behavior and overfitting.
  2. 2.
    Incompatibility with MCMC: The flat density of Uniform priors provides no gradient information to guide sampling algorithms, resulting in inefficient exploration and convergence issues.
Dataset N(0,1) U(-1,1) $\Delta$
Stroke 0.694 0.448 -0.246
Blood 0.583 0.575 -0.008
Heart Failure 0.752 0.724 -0.028
Adult 0.745 0.761 +0.016
Mean 0.694 0.627 -0.066
Table 4: Comparison of Normal vs. Uniform uninformed priors for Bayesian logistic/linear regression.
A.3 Additional experiments
A.3.1 Feature synonyms
We evaluated the impact of feature synonyms on model performance to determine whether changing feature names could introduce bias based on word selection. To do this, we created three sets of synonyms for the features in our evaluation datasets. Table 5 presents examples of these features and their corresponding synonyms. We test three variants: (1) concise, (2) verbose, and (3) clinical feature descriptions. Examples of these variants are provided in Table 5.
Results
Table 6 shows the effect of different grammatical styles for describing features. We find that Feature description style shows modest and inconsistent effects. Across four datasets, we observe an average performance range of 5.7 AUC points between the best and worst description variants. Variant 3 (clinical/formal) achieves the highest mean performance (0.707), outperforming concise (0.679) and verbose (0.664) styles. However, the effect is dataset-dependent: Adult Income shows substantial sensitivity (13.1 point range), with Variant 2 performing poorly (0.636 vs. 0.753/0.767 for other variants), while Stroke exhibits minimal sensitivity (1.9 point range). For medical datasets (Stroke, Heart Failure), clinical terminology appears marginally beneficial, suggesting domain-appropriate phrasing may improve prior extraction. The inconsistency across datasets indicates that while LLMs are not entirely robust to paraphrasing, the sensitivity is unpredictable and relatively small.
Variant Example Prompt
Variant 1 Consider a person who is older. Is this person at risk of stroke? Answer:
Consider a person who has hypertension. Is this person at risk of stroke? Answer:
Consider a person who has fewer months since last donation. Does this person have blood donation likelihood? Answer:
Variant 2 Consider a person who is of older age. Is this person at risk of stroke? Answer:
Consider a person who has a diagnosis of hypertension. Is this person at risk of stroke? Answer:
Consider a person who has donated blood more recently. Does this person have blood donation likelihood? Answer:
Variant 3 Consider a person who exhibits advanced age. Is this person at risk of stroke? Answer:
Consider a person who presents with hypertension. Is this person at risk of stroke? Answer:
Consider a person who demonstrates recent donation activity. Does this person have blood donation likelihood? Answer:
Table 5: Example prompts for feature description variants. All variants use the same template structure but differ in how features are described.
A.3.2 Contextual preambles
We test whether providing domain-knowledge context before feature-effect questions improves prior extraction. For each feature, we prepend an explanatory preamble describing the causal mechanism (e.g., “Stroke risk increases significantly with age due to arterial stiffening and accumulated vascular damage. Based on this medical knowledge,”) before asking the standard question. Table 7 shows example prompts with and without contextual preambles.
Table 6 shows the effect of adding domain-knowledge context before feature-effect questions.
Feature Description Variants Contextual Preambles
Dataset Variant 1 Variant 2 Variant 3 Range $\checkmark$ $\times$ $\Delta$
(Concise) (Verbose) (Clinical) (Baseline)
Stroke 0.685 0.703 0.704 0.019 0.694 0.707 +0.014
Blood Donation 0.567 0.582 0.593 0.026 0.583 0.530 -0.053
Heart Failure 0.713 0.736 0.764 0.051 0.752 0.638 -0.114
Adult Income 0.753 0.636 0.767 0.131 0.745 0.771 +0.026
Mean 0.679 0.664 0.707 0.057 0.693 0.662 -0.032
Table 6: Ablation study results. Left: Effect of feature description style. Right: Effect of contextual preambles.
Condition Example Prompt
No Preamble (Baseline) Consider a person who is older. Is this person at risk of stroke? Answer:
Consider a person who has donated blood more recently. Does this person have blood donation likelihood? Answer:
Consider a person who has higher ejection fraction. Does this person have poor outcome? Answer:
With Preamble Stroke risk increases significantly with age due to arterial stiffening and accumulated vascular damage. Based on this medical knowledge, consider a person who is older. Is this person at risk of stroke? Answer:
People who have donated blood recently are typically more engaged and committed donors. Based on this behavioral pattern, consider a person who has donated blood more recently. Does this person have blood donation likelihood? Answer:
Higher ejection fraction indicates better cardiac pump function and improved prognosis. Based on this cardiac physiology, consider a person who has higher ejection fraction. Does this person have poor outcome? Answer:
Table 7: Example prompts comparing baseline (no preamble) with contextual preamble condition.
Results
Contextual preambles degrade performance on average. Adding domain-knowledge explanations before feature-effect questions reduces mean AUC by 3.2 points (0.693 → 0.662). The effect is highly inconsistent: Adult Income (+2.6 points) and Stroke (+1.4 points) benefit modestly from preambles, while Heart Failure (-11.4 points) and Blood Donation (-5.3 points) suffer substantial degradation. This suggests that explicit reasoning chains interfere with the LLM’s implicit token probability distributions rather than enhancing them. Preambles may bias the model toward generating coherent text completions (which favor the preamble’s stated relationship) rather than revealing the underlying probability mass over answer tokens. For prior extraction via logit probabilities, minimal prompting appears superior to verbose contextual framing.
A.4 Datasets and OOD splits
A.4.1 Public datasets
Adult Income Becker and Kohavi (1996): Predicts whether an individual’s annual income exceeds \$50K using demographic and educational features such as age, work class, education, marital status, occupation, race, sex, and hours worked per week. The binary classification task is based on 48,842 records from the U.S. Census Bureau Adult dataset.
Cervical Cancer Fernandes et al. (2017): Predicts the risk of cervical cancer using demographic factors, sexual health history, and medical test results. Features include age, number of sexual partners, age at first intercourse, smoking habits, hormonal contraceptive use, and HPV test results. The dataset contains 858 patient records collected from clinics in Colombia.
Chronic Kidney Disease Rubini et al. (2015): Classifies the presence of chronic kidney disease using clinical and laboratory measurements, including blood pressure, glucose level, blood urea, serum creatinine, hemoglobin, and albumin. The binary classification task is based on 400 patient records from multiple hospitals.
Bank Marketing Moro et al. (2014): Predicts whether a client will subscribe to a term deposit based on demographic attributes and call-related features such as age, job type, marital status, education, balance, contact communication type, and duration of the last call. The dataset contains 45,211 records collected from Portuguese banking campaigns.
Blood Transfusion Service Center Yeh (2008): Predicts whether a donor will donate blood again using historical donation behavior, including recency, frequency, monetary amount of previous donations, and time since first donation. The dataset consists of 748 donor records from a blood transfusion center in Taiwan.
Stroke Fedesoriano (2020): Predicts the risk of stroke using demographic and health-related features such as age, hypertension, heart disease, glucose level, BMI, smoking status, and gender. The binary classification task is based on 5110 patient records collected from hospitals in India.
Heart Failure UCI Machine Learning Repository (2020): Predicts mortality due to heart failure using clinical measurements including age, ejection fraction, serum creatinine, serum sodium, platelets, and smoking status. The dataset contains 299 patient records collected from hospitals in Portugal.
Diabetes Prediction Mustafatz (2023): Classifies diabetes risk using blood glucose levels, BMI, and other demographic and health-related features. The dataset contains 520 patient records and is used to predict the onset of diabetes in individuals.
Give Me Some Credit Kaggle (2011): Predicts the probability of financial distress within two years using demographic and credit-related features such as age, monthly income, number of past delinquencies, number of open credit lines, and debt-to-income ratio. The dataset consists of 150,000 records from U.S. financial institutions.
Airfoil Self-Noise Brooks et al. (1989): Predicts the noise levels generated by an airfoil using measurements collected by NASA. Features include frequency, angle of attack, chord length, free-stream velocity, and suction side displacement thickness. The regression task is based on 1,503 measurements of airfoil noise under different operating conditions.
Bike Sharing Fanaee-T (2013): Predicts hourly bicycle rental demand based on temporal and weather-related features, including hour of the day, day of the week, temperature, humidity, wind speed, and season. The dataset consists of 17,379 hourly records collected from a bike-sharing system in Washington, D.C.
Combined Cycle Power Plant Tfekci and Kaya (2014): Predicts net hourly electrical energy output using ambient variables such as temperature, pressure, humidity, and exhaust vacuum. The regression task contains 9,568 records collected from a combined cycle power plant over six years.
Concrete Compressive Strength Yeh (1998): Predicts the compressive strength of concrete using mixture components such as cement, blast furnace slag, fly ash, water, superplasticizer, coarse aggregate, and fine aggregate, as well as curing age. The dataset contains 1,030 records of concrete mixtures tested under controlled conditions.
Communities and Crime Redmond (2002): Predicts per-capita violent crime rates in U.S. communities using demographic, socioeconomic, and law enforcement features, including population, income, employment, education, and police presence. The dataset contains 1,994 community records compiled from U.S. government sources.
A.4.2 Private datasets
Head and Neck Cancer Recurrence Dataset
In collaboration with Henry Ford Health, we compiled a binary classification dataset for Head and Neck Cancer Recurrence. This dataset includes comprehensive features about patient demographics, tumor characteristics, treatment regimens, and clinical staging, and predicts the likelihood of cancer recurrence following radiotherapy. The dataset encompasses 508 patients with head and neck cancers, capturing clinical variables including comorbidity indices (Charlson score), smoking history (pack-years), HPV status, primary tumor site localization, and detailed radiation therapy parameters such as treatment duration, prescription dose, and the critical time interval between biopsy and treatment initiation.
This dataset presents a significant class imbalance (390 non-recurrence vs. 118 recurrence cases), reflecting real-world clinical distributions. The dataset is particularly valuable for evaluating model performance under distributional shift, as we created an out-of-distribution test set focusing on elderly patients ($age\geq 75$), representing a clinically meaningful subpopulation that is not present in any LLM pre-training corpora.
Moreover, this dataset highlights why our approach is necessary. Due to its limited sample size, some patterns learned from the data are not medically plausible. Table 8 shows several noticeable relationships between features and the target variable. For instance, the data suggests that marijuana use is associated with a lower risk of Head and Neck Cancer recurrence, which contradicts established medical evidence Kim et al. (2020). This discrepancy is likely due to sampling bias or confounding factors in the small dataset.
LoID priors address this issue by incorporating domain knowledge into the model. In this case, they can enforce a prior that reflects a more realistic relationship between marijuana use and cancer recurrence. As a result, the model is guided away from spurious correlations, making it more reliable and better suited for generalization to real-world populations.
Recurrence Yes No
Sample Size 118 390
CONTINUOUS FEATURES (Mean [Standard Deviation])
Pack Years 41.3 (38.1) 27.3 (28.7)
Biopsy-RT Duration 44.6 (28.2) 46.2 (47.5)
Average Income 82348.7 (32140.4) 90597.7 (36881.5)
CATEGORICAL FEATURES (Mode [%])
Clinical N 2a (50.0) 0 (29.0)
Clinical T 2 (53.7) 2 (57.6)
Radiation Type IMRT Standard (50.4) IMRT Vmat (52.2)
Alcohol Frequent (41.3) No Alcohol Use (45.8)
Smoking History Former Use (45.4) Never Smoker (41.5)
Grade of Differentiation Moderate (68.7) Moderate (79.0)
Neck RT Bilateral (85.1) Bilateral (91.7)
Tobacco Yes (54.5) No (63.2)
Marijuana No (87.6) No (73.5)
First Primary Site = Larynx-supraglottis No (73.6) No (82.0)
First Primary Site = Larynx-glottis No (82.6) No (92.0)
First Primary Site = Oropharynx-base of tongue No (78.5) No (69.5)
First Primary Site = Oropharynx-tonsil No (79.3) No (70.5)
Concurrent Chemotherapy No (90.1) No (95.5)
Table 8: Notable feature–target relationships in the Head and Neck Cancer Recurrence dataset. While some associations, such as those involving tobacco use, pack years, and alcohol consumption, align with clinical expectations, others are misleading. In particular, a larger proportion of marijuana users appear in the non-recurrence group compared to the recurrence group, which may incorrectly suggest a protective effect. A standard logistic regression model may learn this spurious correlation, despite it lacking medical validity.
A.5 Out-of-Distribution splits
Our goal in this research is to evaluate how well a model trained on a small, specific population generalizes to a larger population. To simulate this, we create conditions where only a subset of the full population is available for training, but the aim remains to extract meaningful patterns from the data. Specifically, we first generate train and test sets representing the entire population using a standard 80-20 shuffle split. From the training set, we then select a subset that excludes certain samples entirely, creating a subpopulation that is not fully representative of the overall population. This setup reflects real-world scenarios, such as Head and Neck Cancer Recurrence, where the limited number of training samples can hinder pattern discovery. Table 9 summarizes the split criteria for all datasets.
Binary Classification:
  • Adult (Income Prediction): Training set excludes individuals age <28<28.
  • Chronic Kidney Disease: Training set excludes patients with blood urea <61.75<61.75 mg/dL.
  • Cervical Cancer: Training set excludes women whose age at first sexual intercourse $\leq 15$.
  • Blood Donation: Training set excludes donors with recency >14>14 months (i.e., recent donors are excluded).
  • Stroke: Training set includes only patients age $\leq 32$.
  • Heart Failure: Training set includes only patients with ejection fraction $\geq 30\%$.
  • Diabetes Prediction: Training set includes patients with blood glucose $\leq 126$ mg/dL; testing set includes hyperglycemic patients with glucose $\geq 126$ mg/dL.
  • Bank Marketing: Training set excludes calls with duration >127>127 seconds.
  • Head & Neck Cancer (HNC): Training set includes patients age $\leq 75$.
  • Good Credit Score: Training set excludes individuals age >31>31.
Regression:
  • Airfoil Noise: Training set includes low-frequency conditions (frequency <1600<1600 Hz).
  • Bike Sharing: Training set includes early hours (hour $\leq 6$).
  • Combined Cycle Power: Training set includes moderate temperatures (AT $\leq 13.51^{\circ}$C).
  • Communities & Crime: Training set includes racially diverse areas (racePctWhite $\leq 0.98$).
  • Concrete Strength: Training set includes high-cement mixes (>350>350 kg/m³), where mixture interactions change.
Dataset Type Split Feature OOD Test Condition $P_{\text{IID}}$ $P_{\text{OOD}}$
Binary Classification
Adult Binary Age Age << 28 0.809 0.414
Cervical Cancer Binary First sexual intercourse Age << 15 0.870 0.440
Chronic Kidney Disease Binary Blood urea BU << 61.75 mg/dL 0.993 0.822
Diabetes Prediction Binary Blood glucose Glucose >> 126 mg/dL 0.963 0.883
Stroke Binary Age Age >> 32 0.842 0.608
Heart Failure Binary Ejection fraction EF << 30% 0.849 0.746
Bank Marketing Binary Call duration Duration >> 127 sec 0.878 0.780
Blood Donation Binary Recency Recency >> 14 months 0.748 0.572
Head & Neck Cancer Binary Age Age >> 75 0.998 0.542
Good Credit Score Binary Age Age >> 31 0.872 0.800
Regression
Concrete Strength Regression Cement Cement >> 350 kg/m³ 98.1 469.2
Communities Crime Regression Race (White %) White >> 98% 0.007 0.171
Airfoil Noise Regression Frequency Freq << 1600 Hz 23.7 77.2
Combined Cycle Power Regression Temperature Temp >> 13.51°C 21.2 962.9
Bike Sharing Regression Hour Hour >> 6 19030 77858
Table 9: Out-of-distribution split specifications for all datasets. Train/test splits are created by partitioning on a single feature to induce distributional shift with non-overlapping support.
A.5.1 Rationale
These splits target realistic distribution shifts where:
  1. 1.
    Subpopulation exclusion is plausible: Training data may naturally lack certain demographics (young adults in income surveys), extreme operating conditions (high-frequency turbine tests), or rare clinical presentations (severe heart failure cases).
  2. 2.
    Feature relationships may differ: Age-income correlations differ for young adults vs. middle-aged workers; thermodynamic equations behave differently at temperature extremes.
  3. 3.
    Prior knowledge should help: LLMs may have generalizable knowledge (e.g., "older age increases stroke risk") that transfers to the OOD regime, even if training data provides no direct evidence.
A.6 LoID vs. correlation analysis
In addition to the baselines mentioned in Table 1, we test whether simple correlation analysis might capture similar relationships to LLM priors. To address this, we test a correlation-based baseline that uses feature-target Pearson correlations from the training data as prior means (with fixed $\sigma=1$). Table 10 compares LoID against correlation-based priors. As expected, correlations learned from OOD train sets can mislead the model. In the Bank Marketing dataset where LoID failed because it is dataset-specific and in Cumbine Cycle Power where $\mathcal{N}(0,1)$ could also regain a lot of the performance showing that the train set was actually kinda representing the test set and so correlation studies also got a good AUC. These results further confirm our previous findings.
Binary Classification (AUC) Regression (MSE)
Dataset LoID Corr. $\Delta$ Dataset LoID Corr. $\Delta$
Adult 0.725 0.529 -0.196 Airfoil Noise 60.5 51.4 -9.1
Bank Marketing 0.652 0.770 +0.118 Bike Sharing 78924 69969 -8955
Blood Donation 0.663 0.574 -0.089 Combined Cycle Power 961.9 24.1 -937.8
Cervical Cancer 0.681 0.450 -0.231 Communities & Crime 0.07 0.18 +0.11
Chronic Kidney Disease 0.901 0.575 -0.326 Concrete Strength 121.0 2292.2 +2171.2
Diabetes Prediction 0.906 0.873 -0.033
Heart Failure 0.753 0.742 -0.011
HNC 0.696 0.566 -0.130
Stroke 0.692 0.554 -0.138
Good Credit Score 0.760 0.376 -0.382
Mean 0.741 0.626 -0.142 Mean 16013 14467 -1546
Table 10: LoID (LLM priors) vs. correlation-based priors across 13 datasets (8 classification, 5 regression). For regression, lower MSE is better; negative $\Delta$ indicates correlation outperforms LoID.
A.7 Computational efficiency
We compare the computational complexity and runtime of LoID with AutoElicit and LLMProcesses. For a dataset with $d$ features, the LoID method queries the LLM using $N_{\text{sent}}$ sentence formulations per feature and extracts the probabilities of two tokens (positive and negative). Its time complexity is:
$T_{\text{logit}}=\mathcal{O}(d\times N_{\text{sent}}\times T_{\text{LLM}}),$
where $T_{\text{LLM}}$ denotes the inference time of a single LLM query. In our experiments, $N_{\text{sent}}=10$, resulting in approximately 20 queries per feature.
AutoElicit employs an interactive multi-round elicitation process with iterative refinement, leading to a higher complexity:
$T_{\text{AE}}=\mathcal{O}(d\times N_{\text{rounds}}\times N_{\text{queries}}\times T_{\text{LLM}}),$
where $N_{\text{rounds}}=10$ and $N_{\text{queries}}=10$. This results in more than 28 LLM queries per feature on average.
[…]
LLMProcesses follows a fundamentally different paradigm: it generates predictions for each test instance by sampling from the LLM $N_{\text{samples}}$ times using in-context examples. For a test set of size $n_{\text{test}}$ with $n_{\text{train}}$ training examples in context, its complexity is:
$T_{\text{LLMP}}=\mathcal{O}(n_{\text{test}}\times N_{\text{samples}}\times T_{\text{LLM}}(n_{\text{train}}\cdot d)),$
where $T_{\text{LLM}}(n_{\text{train}}\cdot d)$ reflects the increased inference time caused by long prompts containing $n_{\text{train}}$ in-context examples. Unlike LoID and AutoElicit, which perform prior elicitation once per dataset, LLMProcesses scales linearly with the test set size and incurs substantially longer runtimes (15–20 hours for datasets with 100–400 test instances, compared to minutes for prior elicitation methods).
A.8 Resources and reproducibility
Code availability
To ensure reproducibility, we will release our full codebase, including detailed instructions for running the experiments and reproducing all reported results. We also provide the the posterior files generated by our experiments, as well as the scripts used to generate all figures and tables in the paper.
Data availability
All datasets used in this study except Head and Neck Cancer Recurrence are publicly available through open-access repositories. Further details and references are provided in Section A.4. All preprocessing steps required to obtain the final train/test splits used in our experiments, along with the preprocessed data, are available in our code release.
The Head and Neck Cancer Recurrence dataset will be released upon publication.
Results variability
To assess the stability and determinism of our approach, all experiments were repeated five times. The resulting 95% confidence intervals for AUC were effectively zero (0.000), indicating that our method yields highly consistent results across runs.
Computational resources
  1. 1.
    Infrastructure: Experiments were run on a workstation equipped with 2 NVIDIA RTX 6000 Ada Generation GPUs (each with 50GB VRAM), a Threadripper PRO 7985WX CPU with 64 cores (128 threads) at up to 5.37 GHz, and 256 GB of RAM.
  2. 2.
    Model inference: We used Gemma-2-27B as our main LLM. All inference was performed using both GPUs in parallel. Each feature prompt required about 50–70 seconds to complete across 10 paraphrases, leading to approximately 60 seconds * #features per dataset.
  3. 3.
    Preprocessing: Minimal preprocessing was required and was handled efficiently on a single CPU.
  4. 4.
    Additional resources: Exploratory analyses, including feature selection, were performed using GPT-4o via OpenRouter. Each call took 2–3 seconds and used one CPU thread.